The power delivered by an auxiliary turbine is 500,000 Btu /hr when the steam flow is 4000 lb/hr. Find a. the specific work in ft·lbf/lb b. The power delivered, Hp
1
Expert's answer
2020-03-23T12:22:47-0400

(a)


"\frac{500000}{4000}=125" "\frac{Btu}{lb}" or "125\cdot 778=97250 \frac{ft\cdot lbf}{lb}" Answer


(b)


"1 \frac{Btu}{hr} =0.000393014779 Hp"


"500000\frac{Btu}{hr}=196.5Hp"